# Critical spin models from holographic disorder

### v1

@@ I Introduction
Studying the holographic principle through the lens of quantum information theory has proven to be fruitful for both the field of quantum information and high-energy physics. In particular, within the context of the anti-de Sitter/conformal field theory (AdS/CFT) correspondence Witten (1998); Maldacena (1999), holographic descriptions of quantum information concepts such as entanglement entropy Ryu and Takayanagi (2006); Rangamani and Takayanagi (2017), continuum Almheiri et al. (2015) and tensor-network quantum error correction Pastawski et al. (2015); Jahn and Eisert (2021), and quantum cryptography Apel et al. (2020) have been found. One convenient approach to studying quantum information features of holographic systems is to use finite-dimensional toy models, in particular tensor networks with a hyperbolic lattice geometry Swingle (2012). Such tensor networks are typically constructed as planar $2$-dimensional discretizations of time-slices of AdS2+1 space-time, with open indices at the boundary of a disk-shaped region representing a discretized CFT state. Under a regular discretization of the AdS bulk, these boundary states exhibit quasi-periodic symmetries on the boundary Boyle et al. (2020); Jahn et al. (2020, 2022a); Basteiro et al. (2022, 2023), leading to the question which specific finite-dimensional boundary theories can be described by such tensor network models. The properties of a class of efficiently contractible matchgate tensor networks Valiant (2002); Cai et al. (2007); Bravyi (2009) can be explicitly computed on large hyperbolic tilings, reproducing the average properties of the critical Ising CFT with central charge $c=1/2$ Jahn et al. (2019). In this setting, the boundary Hamiltonian (i.e., the parent Hamiltonian whose ground state coincides with the tensor network state) has been explicitly constructed Jahn et al. (2022b). The resulting nearest-neighbor Hamiltonians obey quasiperiodicity on several length scales and can be well-approximated by the analytical multi-scale quasicrystal ansatz (MQA), which we will describe below in detail. However, the matchgate tensor constraint of these studies restricted the boundary theory to one of free fermions, remote from the interacting boundary theories found in AdS/CFT.
@@
In this work, we overcome this limitation by studying general boundary theories with MQA symmetries beyond its use as a phenomenological model in the hyperbolic tensor network setup of Ref. Jahn et al. (2022b). First, we study the MQA symmetries for the $\{3,7\}$-hyperbolic tiling, describing how these symmetries translate into a sequence of couplings. With these couplings we generate disordered Gaussian Hamiltonians and study their critical behavior. Generalizing this analysis, we also study disordered non-Gaussian Hamiltonians via a matrix product state (MPS) ansatz for the boundary, thus extending beyond the Gaussian regime probed by matchgate tensor networks. We then further generalize to a broad analysis of the MQA for other hyperbolic lattices as well as generic quasiperiodic sequences with and without the MQA. In a nutshell, we investigate the robustness of criticality against quasiperiodic disorder in the following settings:
@@
We find that disordered critical phases with a CFT continuum limit appear generically for those types of the MQA that can be expressed in terms of a “holographically dual” bulk tiling. Conversely, Hamiltonians with non-MQA quasiperiodic symmetries related to hyperbolic tilings, e.g. those symmetries derived from only a single layer of the bulk, are shown to exhibit other critical and non-critical phases without such a continuum limit.
II Results
II.1 Hyperbolic lattices and boundary symmetries
Figure 1: (a): The first three inflation steps of the $\{3,7\}$-hyperbolic lattice in terms of $o$, $a$ and $b$ vertices (gray, blue and orange). Shaded in blue, we track a path along bulk vertices from the center to a given boundary vertex $k$, the product of couplings along which yields the coupling constant $J_{k}$ in the multi-scale quasicrystal ansatz (MQA). (b): (upper panel) Visualization of the MQA for the first three inflation step. The dashed lines denote the bulk path shaded in (a). (lower panel): Visualization of the MQA for the first three inflation steps with symmetrization. (c): We show the coupling constant sequences for various cases, for all of which the disorder strength is the same. (upper panel): The coupling constant sequence $J_{k}$ of the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice after the application of the MQA. (middle panel): The coupling constant sequence $J_{k}$ of the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice without the application of the MQA. The resulting sequence is exactly the two-letter sequence determined by the corresponding inflation rule. (lower panel): The coupling constant sequence $J_{k}$ generated by the function $j_{k}=1+D\cdot\cos(2\pi\frac{1+\sqrt{5}}{2}\cdot k)$ where the parameters have been choosen such that the disorder strength is the same with the two sequences above
@@ II.1 Hyperbolic lattices and boundary symmetries
An AdS time-slice, i.e., the hyperbolic disk, can be symmetrically discretized by a regular hyperbolic tiling characterized by the Schläfli symbols $\{p,q\}$, where $p$ denotes the number of edges of each tile, and $q$ is the number of such tiles that share the same vertex. Hyperbolic triangular tilings are then denoted by $\{3,q\}$ with $q\geq 7$, the $q=7$ case of which is pictured in Fig. 1(a). While the bulk is regular, its boundary is quasiperiodic: Cutting off the tiling after $n$ layers of tiles (starting from a single central tile), the adjacency of open vertices varies along the boundary. Denoting vertices with two, three, and four adjacent edges (up to the cutoff) as $o$, $a$, and $b$, respectively, adding another layer to the tiling is equivalent to applying an inflation rule
@@
to a sequence of letters denoting the previous boundary vertices. In Fig. 1(a), we begin with three $o$ vertices at the center (sequence $ooo$), shaded in gray. Starting from these vertices, we build the lattice step by step, following the inflation rule above, producing a sequence of only $a$ (blue) and $b$ vertices (orange) in the sequence $aaabaaabaaab$. As we repeat this procedure, we produce a sequence of blue and orange vertices that alternates quasiperiodically (in addition to the initial $\mathbb{Z}_{3}$ symmetry). This symmetry also appears in tensor network models built from such geometries Boyle et al. (2020); Jahn et al. (2020, 2022a): Identifying each tile with a $p$-leg tensor and contracting pairs of legs between neighboring tiles/tensors up to the cutoff, one would expect this geometric boundary symmetry to be reflected in the states produced by the tensor network, assuming that the tensors are chosen homogeneously and not break the symmetries of the tiling. In a previous analysis Jahn et al. (2019), the authors used a matchgate tensor network to fill the bulk of the $\{3,7\}$-hyperbolic lattice, respecting its symmetries. This structure leads to boundary states that show a polynomial decay of site-averaged correlations, indicative of critical behavior and following the continuum CFT expectation for the critical Ising model.
Following this work, we now wish to study the critical behavior of boundary theories of such holographic models. In other words, we want to translate the boundary disorder into a sequence of local couplings $\{J_{k}\}$ to build a disordered 1-dimensional Hamiltonian. Written in terms of Majorana modes $\mathrm{\gamma}_{k}$ with $\{\mathrm{\gamma}_{j},\mathrm{\gamma}_{k}\}=2\delta_{j,k}$, one such Hamiltonian is given by
$H_{\text{I}}=\frac{\mathrm{i}}{2}\sum_{k}J_{k}\gamma_{k}\gamma_{k+1}\ ,$ (2)
which is a disordered generalization of the Ising model (for which we would set all odd and even couplings to constants $J/2$ and $h$, respectively). If we set all $J_{k}$ to a constant, $H_{\text{I}}$ would describe the critical point of the Ising model. Generally, we expect arbitrary disorder to interfere with critical behavior. The boundary states of hyperbolic tensor networks, however, do preserve criticality, and they allow for approximate local parent Hamiltonians of the form (2) with varying couplings $J_{k}$ that follow an analytical multi-scale quasicrystal ansatz (MQA) Jahn et al. (2022b, a).
Multi-scale quasicrystal symmetries
The MQA is a one-parameter ansatz class for the couplings $J_{k}$ of (2) that can be formulated for any inflation rule, in particular those corresponding to $\{p,q\}$ hyperbolic tilings. While the MQA reproduces the boundary states of hyperbolic matchgate tensor networks for a particular value of the free parameter, we will consider it in generality.
The MQA is defined from two conditions. The first determines the odd couplings $J_{2k+1}$, corresponding to couplings of Majorana modes within the same physical site, while the even couplings occur between Majorana modes on two neighboring sites. The former are determined from the latter, simply being set to the average of the couplings with adjacent sites:
$J_{2k+1}=\frac{J_{2k}+J_{2k+2}}{2}\ .$ (3)
@@
The second condition of the MQA determines the even couplings $J_{2k}$ from a letter inflation rule. Previous disorder constructions with quasiperiodic symmetry used couplings that alternate between two values $j_{a}$ and $j_{b}$ according to a quasiperiodic two-letter sequence Iglói et al. (2007); Juhász and Zimborás (2007), but the MQA goes on step further: It superimposes the quasiperiodicity from all layers of the inflation steps. This mimics the effect of the hyperbolic tensor network construction, in which each radial layer of tensors acts as a disordered fine-graining transformation. Consider the example of the $\{3,7\}$ tiling in Fig. 1(a): Here we associate “elementary” couplings $j_{a}$ and $j_{b}$ to each vertex type $a$ and $b$, and construct the boundary coupling $J_{k}$ by multiplying all the elementary couplings in previous layers. Keeping track of all vertices deep in the bulk is natural from a holographic perspective, as it is expected that the boundary encodes information at all levels of depth in the bulk, rather than merely that close to the boundary. There is an ambiguity in choosing this path through previous layers, and the naive construction (letter $x$ precedes $y$ if it is included in the inflation rule of $x$) may lead to a sequence that breaks the symmetries of the tiling, as shown in Fig. 1(b). This can be rectified by symmetrizing the inflation rules by splitting the letter $b$, for example $a\to\sqrt{b}aa\sqrt{b}$, where each $\sqrt{b}$ still corresponds to a coupling $j_{b}$. We may also associate a coupling $j_{o}$ with the starting vertices $o$, which appear only once in each sequence and therefore act merely as an overall normalization constant. Similarly, rescaling $j_{a}$ and $j_{b}$ by the same factor only changes the normalization, so that the eigenstates of $H$ only depend on the ratio $r=j_{b}/j_{a}$, resulting in a single free parameter of the $\{3,7\}$ MQA. The matchgate tensor network result of Ref. Jahn et al. (2022b) corresponds to a ratio $r\approx 0.526$.
II.2 Conditions for criticality
Generalizing beyond the non-interacting Ising model, we consider the disordered anti-ferromagnetic Heisenberg model described by the Hamiltonian
@@ II.2 Conditions for criticality
The boundary of the $\{3,7\}$-hyperbolic lattice
One-dimensional uniform spin systems, like the Ising or XXZ-Heisenberg model with constant $J_{k}$ for all sites $k$, show critical behaviour characterized by polynomial correlation decay. This is also the expected behavior for a conformal field theory (CFT), which typically appears as the continuum limit of these spin chain models. Another generic feature of CFT ground states in $1{+}1$ dimensions is a logarithmic scaling of the entanglement entropy $S_{A}=-\mathrm{tr}[\rho_{A}\log\rho_{A}]$ Holzhey et al. (1994). Including finite size corrections to a periodic system of length $L$, the entanglement entropy for a subregion $A$ of length $\ell$ follows the form Calabrese and Cardy (2004, 2009)
@@
Figure 2: (a): Site-averaged entanglement entropy scaling over subsystem size $\ell$ for the 3rd inflation step of the $\{3,7\}$-hyperbolic lattice after applying the MQA. Each colored curve corresponds to a different ratio $r=j_{b}/j_{a}$, thus different disorder. The dashed black curves correspond to the critical behavior of the XX-Heisenberg model, without disorder, in perfect agreement with Eq.(5). (b): Averaged entanglement entropy of $\ell=L/2=29$-spin subsystems over different ratios $r$ (disorder). (c): Averaged 2-point correlation decay over distance $d$, for different ratios $r$. (d): Averaged entanglement entropy scaling over subsystem size $\ell$, for the 3rd inflation step of the $\{3,7\}$-hyperbolic lattice without applying the MQA.
In this work we study disordered spin systems breaking translation invariance, thus requiring site-averaging relevant quantities. Defining $S_{j}(\ell)$ as the entanglement entropy of a compact interval $A=[j,\,(j+\ell-1)\operatorname{mod}L]$, we can write the average over a system with periodic boundary conditions as
@@
For the Gaussian model ($\Delta_{0}=0$) we calculate the entanglement entropy of different subsystems using the covariance matrix formulation Surace and Tagliacozzo (2022) (see Appendix A). This formulation cannot be applied for non-Gaussian cases, such as the XXZ-Heisenberg model ($\Delta_{0}\neq 0$), where tensor network approximations are needed, as we will discuss later.
@@
In Fig. 2(a), we plot $S_{\text{avg}}(\ell)$ for the ground state of a disordered Hamiltonian with coupling constants provided by the 3rd inflation step of the $\{3,7\}$-hyperbolic lattice (58 sites with 116 Majorana modes). The parameter $r=j_{b}/j_{a}$ is the ratio between the values of the blue and orange vertices, which controls the disorder strength: the further the ratio $r$ is from $r_{\mathrm{crit}}=1$, the stronger the disorder. For strong disorder, the systems shows a saturation in the entanglement entropy scaling after a specific length scale, indicative of a gapped system expected to arise in that limit. However, for ratios $r\gtrsim 0.5$ the system shows entanglement entropy scaling in perfect agreement with the uniform, translation-invariant case ($r_{\mathrm{crit}}=1$) following the Calabrese-Cardy ansatz 5 for a CFT. This transition with $r$ becomes apparent in Fig. 2(b), where we plot the averaged entanglement entropy $S_{\text{avg}}(\ell)$ of subsystems of size $\ell=L/2$ for different ratios $r$. There is a distinct plateau around the region $20.5<r<$ 2 where all the disordered models have entanglement entropy values close to the one for the non-disordered model. This plateau indicates that moderate MQA disorder preserves criticality, with the specific ratio $r\sim 0.526$ appearing in matchgate tensor networks in Ref. Jahn et al. (2022b) describing only one point in this critical phase. Interestingly, this point is close to a local maximum in entanglement entropy at $r_{\mathrm{max}}\approx 0.55$, as shown in Fig. 2(b). At this point, $S_{\text{avg}}$ even becomes slightly larger than at the translation-invariant point $r=1$. Note that such a local maximum with the same $S_{\text{avg}}$ also exists around $r\approx 2$, suggesting a symmetry $r\leftrightarrow\frac{1}{r}$ in the critical disordered phase. We can compare the entanglement scaling under MQA disorder to the simpler case of aperiodic disorder with a 2-letter sequence $J_{k}\in(j_{a},j_{b})$ varying according to only a single aperiodic sequence, corresponding to the last layer of Fig. 1(b). In this case, the $r\to 0$ and $r\to\infty$ limits can be described by the strong disorder renormalization group (SDRG), leading to a characteristic $S_{\text{avg}}$ scaling in linear segments with a logarithmic envelope Juhász and Zimborás (2007). Fig. 2(d) shows the gradual transition from the translation-invariant to the disordered case for this simpler type of disorder. Unlike the MQA, this simpler aperiodic disorder does not have a stable critical phase at small disorder; in CFT language, the disorder corresponds to a relevant operator that drives the system to another phase (the critical SDRG phase).
@@
The disordered cases of the free-fermion XX-Heisenberg model show exponential decay for $r<0.5$ and $\ell<L/2$, while the cases for $r\gtrsim 0.5$ show polynomial decay which is close to the uniform critical model. Again, there is a regime of Gaussian disordered cases, derived from the boundary of the hyperbolic tiling by translating it into a sequence of coupling constants, that shows critical behavior. This is also in perfect agreement with the entanglement entropy results above.
@@
Figure 3: Averaged entanglement entropy of subsystems of size $\ell=6$ over different ratios $r$. We consider the ground states of the $\{3,7\}$ MQA-disordered Heisenberg model for $\Delta_{0}=0$ (black line), the non-Gaussian cases for $\Delta_{0}=\{0.2,0.4,0.6,0.8\}$ (orange lines) and the isotropic SU$(2)$ model for $\Delta_{0}=1$ (blue line).

### v2

@@ 1 Introduction
Studying the holographic principle through the lens of quantum information theory has proven to be fruitful for both the field of quantum information and high-energy physics. In particular, within the context of the anti-de Sitter/conformal field theory (AdS/CFT) correspondence [1, 2], holographic descriptions of quantum information concepts such as entanglement entropy [3, 4], continuum [5] and tensor-network quantum error correction [6, 7], and quantum cryptography [8] have been found. One convenient approach to studying quantum information features of holographic systems is to use finite-dimensional toy models, in particular tensor networks with a hyperbolic lattice geometry [9]. Such tensor networks are typically constructed as planar $2$-dimensional discretizations of time-slices of AdS2+1 space-time, with open indices at the boundary of a disk-shaped region representing a discretized CFT state. Under a regular discretization of the AdS bulk, these boundary states exhibit quasi-periodic symmetries on the boundary [10, 11, 12, 13, 14], leading to the question which specific finite-dimensional boundary theories can be described by such tensor network models. The properties of a class of efficiently contractible matchgate tensor networks [15, 16, 17] can be explicitly computed on large hyperbolic tilings, reproducing the average properties of the critical Ising CFT with central charge $c=1/2$ [18]. In this setting, the boundary Hamiltonian (i.e., the parent Hamiltonian whose ground state coincides with the tensor network state) has been explicitly constructed [19]. The resulting nearest-neighbor Hamiltonians obey quasiperiodicity on several length scales and can be well-approximated by the analytical multi-scale quasicrystal ansatz (MQA), which we will describe below in detail. However, the matchgate tensor constraint of these studies restricted the boundary theory to one of free fermions, remote from the interacting boundary theories found in AdS/CFT.
@@
In this work, we overcome this limitation by studying general boundary theories with MQA symmetries beyond its use as a phenomenological model in the hyperbolic tensor network setup of Ref. [19]. First, we study the MQA symmetries for the $\{3,7\}$ hyperbolic tiling, describing how these symmetries translate into a sequence of couplings. With these couplings we generate disordered Gaussian Hamiltonians and study their critical behavior. Generalizing this analysis, we also study disordered non-Gaussian Hamiltonians via a matrix product state (MPS) ansatz for the boundary, thus extending beyond the Gaussian regime probed by matchgate tensor networks. We then further generalize to a broad analysis of the MQA for other hyperbolic lattices as well as generic quasiperiodic sequences with and without the MQA. In a nutshell, we investigate the robustness of criticality against quasiperiodic disorder in the following settings:
@@
We find that disordered critical phases with a CFT continuum limit appear generically for those types of the MQA that can be expressed in terms of a “holographically dual” bulk tiling. Conversely, Hamiltonians with other types of quasiperiodic symmetries related to hyperbolic tilings, e.g. those symmetries derived from only a single layer of the bulk, are shown to exhibit other critical and non-critical phases without such a continuum limit.
2 Results
2.1 Multi-scale quasiperiodic disorder
The multi-scale quasicrystal ansatz (MQA) is a particular class of disordered $(1+1)$-dimensional spin or fermionic chains with nearest-neighbor interactions [19]. It was initially introduced as a disordered version of the transverse-field Ising Hamiltonian
$H_{\text{I}}=\frac{\mathrm{i}}{2}\sum_{k=1}^{2L}J_{k}\,\mathrm{\gamma}_{k}\mathrm{\gamma}_{k+1}\ ,$ (1)
written in terms of $2L$ Majorana modes $\mathrm{\gamma}_{k}$ with $\{\mathrm{\gamma}_{j},\mathrm{\gamma}_{k}\}=2\delta_{j,k}$, and a set $\{J_{k}\}$ of local couplings. The standard translation-invariant form of the Ising model corresponds to setting all odd and even couplings $J_{k}$ to constants $J/2$ and $h$, respectively. In this work we focus on periodic boundary conditions and identify $\mathrm{\gamma}_{2L+1}\equiv-\mathrm{\gamma}_{1}$ for consistency with the usual Ising model definition in the spin picture.
In an MQA, the couplings are defined from two conditions. The first determines the odd couplings $J_{2k+1}$, corresponding to couplings of Majorana modes within the same physical site, while the even couplings occur between Majorana modes on two neighboring sites. The former are determined from the latter, simply being set to the average of the couplings with adjacent sites:
$J_{2k+1}=\frac{J_{2k}+J_{2k+2}}{2}\ .$ (2)
The second condition of the MQA determines the even couplings $J_{2k}$ from a letter inflation rule. Previous disorder constructions with quasiperiodic symmetry used couplings that alternate between two values $j_{a}$ and $j_{b}$ according to a quasiperiodic two-letter sequence [20, 21], which is scaled up to larger sequences in each inflation step by replacing single letters with longer subsequences. For example, the quasiperiodic Fibonacci sequence emerges from iteratively applying the inflation rule $a\mapsto ab$, $b\mapsto a$. The MQA, however, goes one step further: It superimposes the quasiperiodicity from all steps of such an inflation rule by multiplying the couplings $j_{a},j_{b}$ associated with each letter across iteration steps to form the physical coupling constants $J_{k}$. This mimics a critical renormalization-group transformation in which self-similar disorder is introduced at every length scale, as in holographic tensor network constructions [10, 11, 12].
@@ 2.1 Multi-scale quasiperiodic disorder
Letter inflation rules appear naturally when describing layers of regular hyperbolic tilings, i.e., tessellations of the hyperbolic disk by regular $p$-gons. These can be classified by the Schläfli symbols $\{p,q\}$, where $p$ denotes the number of edges of each tile, and $q$ is the number of such tiles that share the same vertex. Hyperbolic triangular tilings are then denoted by $\{3,q\}$ with $q\geq 7$, the $q=7$ case of which is pictured in Fig. 1(a). While the bulk is regular, its boundary is quasiperiodic: Cutting off the tiling after $n$ layers of tiles (starting from a single central tile), the adjacency of open vertices varies along the boundary. Denoting vertices with two, three, and four adjacent edges (up to the cutoff) as $o$, $a$, and $b$, respectively, adding another layer to the tiling is equivalent to applying the inflation rule
@@
to a sequence of letters denoting the previous boundary vertices. In Fig. 1(a), we begin with three $o$ vertices at the center (sequence $ooo$), shaded in gray. Starting from these vertices, we build the lattice step by step, following the inflation rule above, producing a sequence of only $a$ (blue) and $b$ vertices (orange) in the sequence $aaabaaabaaab$. As we repeat this procedure, we produce a sequence of blue and orange vertices that alternates quasiperiodically (in addition to the initial $\mathbb{Z}_{3}$ symmetry). This symmetry also appears in tensor network models built from such geometries [10, 11, 12]: Identifying each tile with a $p$-leg tensor and contracting pairs of legs between neighboring tiles/tensors up to the cutoff, one would expect this geometric boundary symmetry to be reflected in the states produced by the tensor network, assuming that the tensors are chosen homogeneously and not break the symmetries of the tiling. This assumption includes random tensor networks on such geometries [22], where sample averaging will smooth out any deviations from the lattice symmetries, as well as holographic codes such as those introduced in Ref. [6], whose code states also exhibit MQA symmetries [12]. A previous analysis utilized matchgate tensor networks, based on tensors with an efficiently computable free-fermion representation, to fill the bulk of the $\{3,7\}$ hyperbolic lattice, respecting its symmetries [18]. This structure led to boundary states that show a polynomial decay of site-averaged correlations, indicative of critical behavior and following the continuum CFT expectation for the critical Ising model. Generally, we expect arbitrary disorder to interfere with critical behavior; the boundary states of hyperbolic tensor networks, however, do preserve criticality, and they allow for approximate local parent Hamiltonians of the form (1) with varying couplings $J_{k}$ that follow an MQA [19, 12].
@@
Consider the example of the $\{3,7\}$ hyperbolic tiling in Fig. 1(a): Here we associate “elementary” couplings $j_{a}$ and $j_{b}$ to each vertex type $a$ and $b$, and construct the physical couplings $J_{k}$ in Eq. (1), associated with modes on the tiling boundary, by multiplying all the elementary couplings across the bulk layers. Keeping track of all vertices deep in the bulk is natural from a holographic perspective, as it is expected that the boundary encodes information at all levels of depth in the bulk, rather than merely those that are close to the boundary. There is an ambiguity in choosing this path through previous layers, and the naive construction (letter $x$ precedes $y$ if it is included in the inflation rule of $x$) may lead to a sequence that breaks the symmetries of the tiling, as shown in Fig. 1(b). This can be rectified by symmetrizing the inflation rules by splitting the letter $b$, for example $a\to\sqrt{b}aa\sqrt{b}$, where each $\sqrt{b}$ still corresponds to a coupling $j_{b}$. We may also associate a coupling $j_{o}$ with the starting vertices $o$, which appears only once in each sequence and therefore acts merely as an overall normalization constant. Similarly, rescaling $j_{a}$ and $j_{b}$ by the same factor only changes the normalization, so that the eigenstates of $H$ only depend on the ratio $r=j_{b}/j_{a}$, resulting in a single free parameter of the $\{3,7\}$ MQA. In the matchgate tensor network result of Ref. [19], the $\{3,7\}$ tiling led to a ratio $r\approx 0.526$ depending on the geometric irregularity of the tiling boundary at a finite cutoff. Tilings with higher curvature, i.e., $\{p,q\}$ tilings with larger $p$ and $q$, exhibit a more irregular boundary and hence more disorder for the same value of $r$, so they are expected to exhibit disordered Ising-like MQA phases at values of $r$ closer to $1$.
Figure 1: (a): The first three inflation steps of the $\{3,7\}$ hyperbolic lattice in terms of $o$, $a$ and $b$ vertices (gray, blue and orange). Shaded in blue, we track a path along bulk vertices from the center to a given boundary vertex $k$, the product of couplings along which yields the coupling constant $J_{k}$ in the multi-scale quasicrystal ansatz (MQA). (b): (upper panel) Visualization of the MQA for the first three inflation step. The dashed lines denote the bulk path shaded in (a). (lower panel) Visualization of the MQA for the first three inflation steps with symmetrization. (c): We show the coupling constant sequences for various cases, for all of which the disorder strength is the same. (upper panel) The coupling constant sequence $J_{k}$ of the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice after the application of the MQA. (middle panel) The coupling constant sequence $J_{k}$ of the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice without the application of the MQA. The resulting sequence is exactly the two-letter sequence determined by the corresponding inflation rule. (lower panel) The coupling constant sequence $J_{k}$ generated by the quasiperiodic function $j_{k}=1+D\cdot\cos(2\pi\frac{1+\sqrt{5}}{2}\cdot k)$ where the parameters have been choosen such that the disorder strength is the same as for the two sequences above.
Though the MQA was previously applied to a disordered Ising model (1), which is non-interacting and therefore exactly solvable with Gaussian techniques, it can also be generalized to other disordered Hamiltonians with nearest-neighbor couplings. In particular, we consider the disordered anti-ferromagnetic Heisenberg model described by the Hamiltonian
@@ 2.2 Criticality in Gaussian MQA
One-dimensional uniform spin systems, like the Ising or XXZ-Heisenberg model with constant $J_{k}$ for all sites $k$, show critical behaviour characterized by polynomial correlation decay. This is also the expected behavior for a conformal field theory (CFT), which typically appears as the continuum limit of these spin chain models. Another generic feature of CFT ground states in $1{+}1$ dimensions is a logarithmic scaling of the entanglement entropy $S_{A}=-\mathrm{tr}[\rho_{A}\log\rho_{A}]$ [23]. Including finite size corrections to a periodic system of length $L$, the entanglement entropy for a subregion $A$ of length $\ell$ follows the form [24, 25]
@@
Figure 2: (a): Site-averaged entanglement entropy scaling over subsystem size $\ell$ for the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice after applying the MQA. Each colored curve corresponds to a different ratio $r=j_{b}/j_{a}$, thus different disorder. The dashed black curves correspond to the critical behavior of the XX-Heisenberg model, without disorder, in perfect agreement with Eq. (5). (b): Averaged 2-point correlation decay over distance $d$, for different ratios $r$. (c): Averaged entanglement entropy of $\ell=L/2=29$-spin subsystems over different ratios $r$ (disorder). (d): Averaged entanglement entropy scaling over subsystem size $\ell$, for the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice without applying the MQA. (e): Averaged entanglement entropy scaling over subsystem size $\ell$, for the 4th inflation step of the $\{3,7\}$ hyperbolic lattice without applying the MQA. (f): Averaged entanglement entropy of $\ell=29$-spin subsystems over different ratios $r$ for the 3rd, 4th, 5th and 6th inflation step of the $\{3,7\}$ hyperbolic lattice without applying the MQA.
Here we study disordered spin systems breaking translation invariance, thus requiring site-averaging relevant quantities. Defining $S_{j}(\ell)$ as the entanglement entropy of a compact interval $A=[j,\,(j+\ell-1)\operatorname{mod}L]$, we can write the average over a system with periodic boundary conditions as
@@
For the Gaussian model ($\Delta_{0}=0$) we calculate the entanglement entropy of different subsystems using the covariance matrix formulation [26] (see Appendix A). This formulation cannot be applied for non-Gaussian cases, such as the XXZ-Heisenberg model ($\Delta_{0}\neq 0$), where tensor network approximations are needed, as we will discuss later.
@@
In Fig. 2(a), we plot $S_{\text{avg}}(\ell)$ for the ground state of a disordered Hamiltonian with coupling constants provided by the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice (58 sites with 116 Majorana modes). The parameter $r=j_{b}/j_{a}$ is the ratio between the values of the blue and orange vertices, which controls the disorder strength: The further the ratio $r$ is from $r_{\mathrm{crit}}=1$, the stronger the disorder. For strong disorder, the systems shows a saturation in the entanglement entropy scaling after a specific length scale, indicative of a gapped system expected to arise in that limit. However, for ratios $r\gtrsim 0.5$ the system shows entanglement entropy scaling in perfect agreement with the uniform, translation-invariant case ($r_{\mathrm{crit}}=1$) following the Calabrese-Cardy ansatz (5) for a CFT. This transition with $r$ becomes apparent in Fig. 2(c), where we plot the averaged entanglement entropy $S_{\text{avg}}(\ell)$ of subsystems of size $\ell=L/2$ for different ratios $r$. There is a distinct plateau around the region $20.5<r<$ 2 where all the disordered models have entanglement entropy values close to the one for the non-disordered model. This plateau indicates that moderate MQA disorder preserves criticality, with the specific ratio $r\sim 0.526$ appearing in matchgate tensor networks in Ref. [19] describing only one point in this critical phase. Interestingly, this point is close to a local maximum in entanglement entropy at $r_{\mathrm{max}}\approx 0.55$, as shown in Fig. 2(b). At this point, $S_{\text{avg}}$ even becomes slightly larger than at the translation-invariant point $r=1$. Note that such a local maximum with the same $S_{\text{avg}}$ also exists around $r\approx 2$, suggesting a symmetry $r\leftrightarrow\frac{1}{r}$ in the critical disordered phase. The appearance of such an approximate symmetry at small $|r-1|$ is likely a result of a dependence of critical behavior on the overall disorder strength, which exhibits a similar symmetry; for example, if we measure disorder strength in terms of the standard deviation of the sequence of couplings $J_{k}$, we find the same approximate $r\leftrightarrow\frac{1}{r}$ symmetry. We can compare the entanglement scaling under MQA disorder to the case of aperiodic disorder with a 2-letter sequence $J_{k}\in(j_{a},j_{b})$ taken from the letters at only the last inflation step that generated a corresponding MQA (e.g. the last row of Fig. 1(b)). In our analysis, we will use the shorthand non-MQA to refer to this simpler type of quasiperiodic disorder. In this case, the $r\to 0$ and $r\to\infty$ limits can be described by the strong disorder renormalization group (SDRG), as local pairs of sites coupled by either $j_{a}\gg j_{b}$ or $j_{b}\gg j_{a}$ can be iteratively replaced by singlets. This singlet-based renormalization process leads to a characteristic $S_{\text{avg}}$ scaling in linear segments with a logarithmic envelope [21]. Figs. 2(d) and (e) show the gradual transition from the translation-invariant to the disordered case for this simpler type of disorder, at $n=3$ and $n=4$ inflation steps. Note that the logarithmic envelope of the entanglement entropy for small subsystem sizes $\ell$ has a smaller prefactor than the translation-invariant value of $1/3$, showing a transition to a different phase in the large-disorder limit (this deviation becomes increasingly visible at larger $n$). Unlike the MQA, this simpler aperiodic disorder therefore does not have a stable critical phase at small disorder: As we show in Fig. 2(f), any deviation from the translation-invariant case $r=1$ results in an immediate decline of subsystem entanglement entropy. In CFT language, this disorder corresponds to a relevant operator that drives the system to another phase (the critical SDRG phase).
@@
For $\Delta_{0}=0$, the Hamiltonian of the Heisenberg model (4) becomes Gaussian, thus of the form of (1), so we can calculate the correlation decay as described above. The disordered cases of the free-fermion XX-Heisenberg model show exponential decay for $r<0.5$ and $\ell<L/2$, while the cases for $r\gtrsim 0.5$ show polynomial decay which is close to the uniform critical model, as seen in Fig. 2(b). Again, there is a regime of Gaussian disordered cases, derived from the boundary of the hyperbolic tiling by translating it into a sequence of coupling constants, that shows critical behavior. This is also in perfect agreement with the entanglement entropy results above.
@@
Figure 3: (a): Averaged entanglement entropy of subsystems of size $\ell=6$ over different ratios $r$. We consider the ground states of the $\{3,7\}$ MQA-disordered Heisenberg model (3rd inflation step) for $\Delta_{0}=0$ (black line), the non-Gaussian cases for $\Delta_{0}=\{0.2,0.4,0.6,0.8\}$ (orange lines) and the isotropic SU$(2)$ model for $\Delta_{0}=1$ (blue line). (b): Spin-spin correlations decay of the ground state of the $\{3,7\}$ MQA-disordered Heisenberg model for $\Delta_{0}=0.5$. (c): Spin-spin correlations decay of the ground state of the $\{3,7\}$ MQA-disordered Heisenberg model for $\Delta_{0}=1$ (SU$(2)$-symmetric XXZ-model).
2.3 Criticality in Non-Gaussian MQA